A mouse is moving in a straight line at a steady speed of 2 m/sec for 10 seconds. How far (in meters) did the mouse travel?
Answer by jim_thompson5910(35256)

(2 m/sec)*(10 sec) = (2*10)*(m/sec * sec) = 20 meters

Notice how the "sec" units will cancel